Circular No. 7908

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dc D. Magee, R. Bouwens, and G. Illingworth, University of California,

Santa Cruz; H. Ford, N. Benitez, J. Blakeslee, N. Cross, C. Gronwall, and Z. Tsvetanov, John Hopkins University; M. Clampin and G. Hartig, Space Telescope Science Institute; and the Advanced Camera for Surveys (ACS) Science Team report the discovery of an apparent supernova (m_{AB} = 22.70 ± 0.15; cf. Oke 1974, Ap.J. Suppl. 27, 21) in the Hubble Deep Field North (HDFN), detected in exposures obtained with the Hubble Space Telescope’s ACS (+ F775W filter) on May 11.39 UT. SN 2002dc is located 0′′.8 east of the center of the galaxy 2-264.1 [as designated by Williams et al. (1996, A.J. 112, 1335), who give its position as α = 12^{h}36^{m}49^{s}.756, δ = +62^{o}13′13′′.09 (equinox J2000.0); Cohen et al. (1996, Ap.J. 471, L5) reported its redshift as 0.475]. The new object is not present in HST WFPC2 F814W images of the HDFN taken in Dec. 1995 (10σ limiting mag m_{AB} = 27.60, determined by Williams et al.). Tsvetanov, Blakeslee, Ford, Magee, Bouwens, Illingworth; A. Riess, Space Telescope Science Institute; and the ACS Science Team further report that SN 2002dc was detected in ACS HDFN exposures (+ F850LP filter), at m_{AB} = 23.15 ± 0.15 (photometric calibration not yet completed), on May 21.17 UT and in the G800L grism on May 11.52; the grism spectrum is consistent with that of a type-I supernova (and most likely a type-Ia event) at this redshift, a month or more after maximum and having well-developed nebular emission.

SAX J1819.3–2525 D. Campbell-Wilson and R.W. Hunstead, University of Sydney, report

the detection with the Molonglo Observatory Synthesis Telescope of a radio counterpart of the x-ray transient SAX J1819.3–2525 = V4641 Sgr (IAUC 7906). Short observations on May 21.89 and 22.80 UT (mid-exposure) yielded 843-MHz flux densities of 32 ± 5 and 133 ± 4 mJy, respectively, coincident with the optical position (IAUC 7277).

SUPERNOVA 2002db IN NGC 5683 S. Nakano, Sumoto, Japan, reports an independent discovery of SN

2002db (cf. IAUC 7906) by R. Kushida at Yatsugatake South Base Observatory, who notes unfiltered mag 16.9 and the following position end figures from a CCD frame taken on May 24.729 UT: 51^{s}.66, 53′′.3.

2002 May 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green